Every bacteria cell divides in 1 min and it take half an hour to fill the cup. How much time it will take to fill half the cup?

Grade:12

1 Answers

Vasanth SR
6 years ago
The question is incomplete and the answer can’t be processed.

A Bacterial cell divides once every minute. It takes one hour to fill a cup. How much time will it take to fill half the cup?
(1) 59 minute
(2) 30 minute
(3) 29 minute
(4) 20 minute

Solution:

b(n) = 2^n

b(60) = 2^60 = 1 cup

1/4 cup = 2^60/4 = 2^58

Time = n = 58 minutes